# Low-resource domain adaptation while minimizing energy and hardware resource consumption

### v1

@@ Abstract
Training Large Language Models (LLMs) is costly in terms of energy, hardware, and annotated data, often resulting in a positionality rooted in predominant cultures and values Santy et al. (2023). Domain adaptation has emerged as a promising strategy to better align models with diverse cultural and value contexts Hershcovich et al. (2022), but its computational cost remains a significant barrier, particularly for research groups lacking access to large-scale infrastructure. In this paper, we evaluate how the use of different numerical precisions and data parallelization strategies impacts both training speed—as a proxy to energy and hardware consumption—and model accuracy, with the goal of facilitating domain adaptation in low-resource environments. Our findings are relevant to any setting where energy efficiency, accessibility, or limited hardware availability are key concerns.
@@ 1 Introduction
In this work, we explore how to make this adaptation process more efficient and accessible in low-resource settings. We focus on adapting the BERTBase model to the VizWiz-VQA Gurari et al. (2018) dataset, which contains images paired with visual questions asked by people with visual impairments during their daily activities. Figure 1 presents representative samples from this dataset, highlighting key challenges such as blurring, rotated text, and subjective queries that can severely impact model performance.
@@ 3 Methods
To assess the model’s domain adaptation performance across multi-GPU setups, we designed experiments involving various parallelization strategies, precision settings, and parameter tuning approaches.
@@ 3.2 Training process
We evaluated two precision modes: standard 32-bit floating-point (FP32) and 16-bit floating-point (FP16) using Automatic Mixed Precision (AMP). Since BF16 is not supported across all GPU architectures, and that TF32 is limited to specific NVIDIA hardware, we prioritized approaches that are widely compatible with diverse GPU platforms and training frameworks such as PyTorch and TensorFlow Abadi et al. (2016), aligning with our goal of exploring strategies accessible to resource-constrained institutions.
@@ 4 Experiments and results
We conducted an evaluation of the optimization strategies applied during the domain adaptation process, focusing on the impact of batch size, parallelization strategies, and numerical precision modes. The analysis covers key metrics such as training performance—measured in $e$pochs $p$rocessed $p$er $m$inute (eppm)—alongside memory usage, power consumption, and model convergence. This assessment aims to clarify the trade-offs between efficiency gains and model performance across different training configurations.
@@ 4.2 Parallelization strategies
We compared baseline measurements obtained from domain adaptation using a single-GPU against the performance of multi-GPU setups with the DP and DDP strategies, using both FP32 and AMP-FP16 precision. With FP32 (see Figure 3(a)), DDP consistently outperformed DP across all tested batch sizes, achieving up to 2× the performance observed with a single GPU. Both parallelization strategies exhibited notable improvements over single-GPU training starting from a batch size of 64. However, for smaller batch sizes, synchronization and communication overhead between GPUs likely offset performance gains, making DP the least favorable option in these cases.
@@ 4.3 Memory usage analysis
Figure 4: GPU memory usage comparison between Data Parallel (DP, top) and Distributed Data Parallel (DDP, bottom). Both plots compare memory consumption across different batch sizes and numerical precision modes (FP32 vs. AMP-FP16). This highlights a clear trade-off between memory usage and training speed.
@@ 4.4 Power usage and consumption
Figure 5: GPU utilization and power consumption during domain adaptation on a single GPU, comparing different batch sizes and numerical precisions (FP32 vs. AMP-FP16). The top row shows GPU utilization over time, while the bottom row displays power consumption (in Watts). The results highlighting the reduction in both GPU load and energy consumption when using AMP-FP16.
@@ 4.5 Model convergence and accuracy
Beyond efficiency metrics, we evaluated whether the proposed optimization strategies had any impact on model performance. To this end, we tracked masked word prediction accuracy on the validation set using the DDP strategy, which consistently delivered the best performance in previous tests. We compared results between single-GPU and dual-GPU setups, evaluating both FP32 and AMP-FP16 precisions. All tests were conducted using the same batch size configurations as in earlier experiments, with learning rates properly scaled. As shown in Figure 7, no significant differences in final model accuracy were observed across precision modes or hardware configurations. These findings suggest that using lower numerical precision is indeed feasible without compromising training quality.
Figure 7: Validation accuracy curves across training epochs using the Distributed Data Parallel (DDP) strategy under both FP32 and AMP-FP16 precision modes. The top plots correspond to single-GPU training, while the bottom plots show results with a dual-GPU setup. No significant differences in final accuracy were observed across precision modes or hardware configurations.
@@ 5 Conclusion
This work systematically investigated optimization strategies for enhancing the efficiency of BERT domain adaptation on GPU infrastructures. We compared Data Parallel (DP) and Distributed Data Parallel (DDP) training under varying batch sizes and precision modes (FP32 and AMP-FP16).
@@
Our analysis focused on FP32 and AMP-FP16, as they are widely supported and commonly used in full-model training. While lower-bit formats like INT8 or INT4 offer appealing efficiency gains, they are typically optimized for inference or adapter-based fine-tuning, which fall outside the scope of this work. Future research could explore intermediate precision formats such as BF16 and TF32, which promise further acceleration without compromising training stability.

### v2

@@ Abstract
Training Large Language Models (LLMs) is costly in terms of energy, hardware, and annotated data, often resulting in a positionality rooted in predominant cultures and values Santy et al. (2023). Domain adaptation has emerged as a promising strategy to better align models with diverse cultural and value contexts Hershcovich et al. (2022), but its computational cost remains a significant barrier, particularly for research groups lacking access to large-scale infrastructure. In this paper, we evaluate how the use of different numerical precision formats and data parallelization strategies impacts both training speed—as a proxy to energy and hardware consumption—and model accuracy, with the goal of facilitating domain adaptation in low-resource environments. Our findings are relevant to any setting where energy efficiency, accessibility, or limited hardware availability are key concerns.
@@ 1 Introduction
In this work, we explore how to make this adaptation process more efficient and accessible in low-resource settings. We focus on adapting the BERTBase [^1] https://huggingface.co/google-bert/bert-base-uncased model to the VizWiz-VQA Gurari et al. (2018) dataset, which contains images paired with visual questions asked by people with visual impairments during their daily activities. Figure 1 presents representative samples from this dataset, highlighting key challenges such as blurring, rotated text, and subjective queries that can severely impact model performance.
@@ 3 Methods
To assess the model’s domain adaptation performance across multi-GPU setups, we designed experiments involving different parallelization strategies, numerical precision formats, and parameter tuning approaches.
@@ 3.2 Training process
We evaluated two numerical precision formats: standard 32-bit floating-point (FP32) and 16-bit floating-point (FP16) using Automatic Mixed Precision (AMP). Since BF16 is not supported across all GPU architectures, and that TF32 is limited to specific NVIDIA hardware, we prioritized approaches that are widely compatible with diverse GPU platforms and training frameworks such as PyTorch and TensorFlow Abadi et al. (2016), aligning with our goal of exploring strategies accessible to resource-constrained institutions.
@@ 4 Experiments and results
We conducted an evaluation of the optimization strategies applied during the domain adaptation process, focusing on the impact of batch size, parallelization strategies, and numerical precision formats. The analysis covers key metrics such as training performance—measured in $e$pochs $p$rocessed $p$er $m$inute (eppm)—alongside memory usage, power consumption, and model convergence. This assessment aims to clarify the trade-offs between efficiency gains and model performance across different training configurations.
@@ 4.2 Parallelization strategies
We compared baseline measurements obtained from domain adaptation using a single-GPU against the performance of multi-GPU setups with the DP and DDP strategies, under both FP32 and AMP-FP16 numerical precision formats. With FP32 (see Figure 3(a)), DDP consistently outperformed DP across all tested batch sizes, achieving up to 2× the performance observed with a single GPU. Both parallelization strategies exhibited notable improvements over single-GPU training starting from a batch size of 64. However, for smaller batch sizes, synchronization and communication overhead between GPUs likely offset performance gains, making DP the least favorable option in these cases.
@@ 4.3 Memory usage analysis
Figure 4: GPU memory usage comparison between Data Parallel (DP, top) and Distributed Data Parallel (DDP, bottom). Both plots compare memory consumption across different batch sizes and numerical precision formats (FP32 vs. AMP-FP16). This highlights a clear trade-off between memory usage and training speed.
@@ 4.4 Power usage and consumption
Figure 5: GPU utilization and power consumption during domain adaptation on a single GPU, comparing different batch sizes and numerical precision formats (FP32 vs. AMP-FP16). The top row shows GPU utilization over time, while the bottom row displays power consumption (in Watts). The results highlighting the reduction in both GPU load and energy consumption when using AMP-FP16.
@@ 4.5 Model convergence and accuracy
Beyond efficiency metrics, we evaluated whether the proposed optimization strategies had any impact on model performance. To this end, we tracked masked word prediction accuracy on the validation set using the DDP strategy, which consistently delivered the best performance in previous tests. We compared results between single-GPU and dual-GPU setups, evaluating both FP32 and AMP-FP16 precisions. All tests were conducted using the same batch size configurations as in earlier experiments, with learning rates properly scaled. As shown in Figure 7, no significant differences in final model accuracy were observed across the numerical precision formats or hardware configurations used. These findings suggest that using lower numerical precision is indeed feasible without compromising training quality.
Figure 7: Validation accuracy curves across training epochs using the Distributed Data Parallel (DDP) strategy under both FP32 and AMP-FP16 numerical precision formats. The top plots correspond to single-GPU training, while the bottom plots show results with a dual-GPU setup. No significant differences in final accuracy were observed across the numerical precision formats or hardware configurations used.
@@ 5 Conclusion
In this study, we conduct a systematic evaluation of optimization strategies aimed at improving the efficiency of BERT domain adaptation on GPU infrastructures. We compare Data Parallel (DP) and Distributed Data Parallel (DDP) strategies during the adaptation process across a range of batch sizes and numerical precision formats (FP32 and AMP-FP16).
@@
Our analysis focused on FP32 and AMP-FP16, as they are widely supported and commonly used in full-model training. While lower-bit formats like INT8 or INT4 offer appealing efficiency gains, they are typically optimized for inference or adapter-based fine-tuning, which fall outside the scope of this work. Future research could explore intermediate numerical precision formats such as BF16 and TF32, which promise further acceleration without compromising training stability.